Logiweb bibliography of my factorial

Up Help

[0] my factorial

[1] base

The pyk compiler, version 0.grue.20050603 by Klaus Grue GRD-2005-06-16.UTC:17:53:35.048381 = MJD-53537.TAI:17:54:07.048381 = LGT-4625661247048381e-6